# Robustness and Cybersecurity in the EU Artificial Intelligence Act

### v1

@@ 5. Requirements for High-Risk AI Systems
In addition to these attacks, Art. 15(5)(iii) AIA lists ’model flaws’ as an AI-specific vulnerability. This is a vague legal term and lacks an established counterpart in the ML literature. In software contexts, the word flaw often refers to so-called bugs, which are typically the result of human errors in the coding process (Kumar and Anderson, 2023; Nissenbaum, 1996). However, the term model flaw follows the list of attacks outlined above, which are instead designed to exploit the default properties of a properly functioning ML model, and are not directly the results of errors in the coding process. Thus, it is unclear what model flaw refers to in this context, and whether technical solutions are only expected to address traditional bugs or coding errors, or whether they should address other ways of exploiting AI-specific vulnerabilities. Given that the term is situated within the cybersecurity requirements for AI system outlined in Art. 15(5) AIA, we argue that the term model flaws should be interpreted as flaws that enable the exploitation of AI-specific vulnerabilities. Technical standards and guidelines by the EU Commission should define model flaws more clearly and provide guidelines for technical solutions to address these model flaws. This should take into account the arms race between attacker and defender in the realm of adversarial robustness, where both parties are continuously adapting their strategies to outmaneuver the other (Chen et al., 2017b). This makes it infeasible to anticipate and counter all potential attacks that target AI-specific vulnerabilities.
@@ 6. Requirements for General-Purpose AI Models With Systemic Risk
In the previous section, we examined HRAIS requirements. To further elucidate them, we study GPAIMs with systemic risk, highlighting similarities and differences. The AIA establishes legal requirements for GPAIMs, such as multimodal large language models (OpenAI, 2023; Gemini Team et al., 2023), which can perform tasks beyond their original training objective (Gutierrez et al., 2023). GPAIM can be standalone or embedded in an HRAIS, with the latter requiring compliance with both GPAIM and HRAIS requirements. The AIA distinguishes between GPAIM with systemic risks and those without. Art. 3(65) AIA defines ‘systemic risk’ as the risk that is specific to the high-impact capabilities of GPAIMs that have a “significant impact” on the market, public health, safety, security, fundamental rights, or society. [^19] A systemic risk is presumed when the cumulative computation during training exceeds $10^{25}$ Floating-Point Operations Per Second (FLOPS). GPAIMs with fewer FLOPS may still be classified as posing a systemic risk under Art. 51(1) AIA. There is an ongoing debate over this threshold and if a model’s complexity truly reflects its risk level (Hacker, 2024; Novelli et al., 2024; Pehlivan, 2024; Kutterer, 2023). Thresholds and criteria can be modified by the EU Commission, see Appendix B GPAIMs without systemic risks are exempt from robustness and cybersecurity obligations (Art. 53 AIA ff.).
@@
Art. 55(1)(d) AIA mandates “an adequate level of cybersecurity protection” for GPAIMs with systemic risk. Rec. (115) further details this cybersecurity requirement. It mandates cybersecurity protection against “malicious use or attacks” and lists specific adversarial threats, such as “accidental model leakage, unauthorised releases, circumvention of safety measures”, “cyberattacks”, or “model theft”. Notably, several of these threats have direct counterparts in the ML literature on adversarial robustness and privacy for large generative models, such as the circumvention of safety measures (jailbreaking) or model theft (Yao et al., 2024; Li et al., 2023; Wang et al., 2023). Although Art. 55(1)(d) AIA does not define the term ‘cyberattacks’, we infer that it includes the attacks exploiting AI-specific vulnerabilities mentioned in Art. 15 AIA (see Section 5.3). These attacks are studied in the field of adversarial robustness and can also affect GPAIMs (Qiang et al., 2024; Das et al., 2024; Yan et al., 2024; Schwinn et al., 2024; Vitorino et al., 2024)—even though specific ML techniques may be necessary to address GPAIM-specific challenges. This relation underscores that the concepts and problems explored under adversarial robustness are reflected in the term ’cybersecurity’ as used in Art. 55(1)(d) AIA. To ensure the cybersecurity of a GPAIMs with systemic risk, providers must conduct and document internal and/or external adversarial testing of the model, such as red teaming. [^20] In this context, red teaming refers to stress testing AI models by simulating adversarial attacks (Feffer et al., 2024), such as linguistic or semantic attacks against LLMs (Shi et al., 2024), whereas traditional cybersecurity red teaming focuses on assessing entire systems or networks (Teichmann and Boticiu, 2023).
While the AIA mandates robustness requirements for HRAIS, we observe that it does not impose an explicit equivalent legal requirement for GPAIMs, regardless of whether they present a systemic risk or not. Specifically, neither Art. 55 AIA nor Rec. (155) address unintentional causes for deviations from consistent performance. In Section 5.2, we stated that non-adversarial robustness is reflected in the term robustness in Art. 15 AIA. Consequently, GPAIMs, which are not required to fulfill any robustness requirement, are not mandated to be resilient against performance issues, such as data distribution shifts or noisy data. The AIA itself does not provide an explanation for the omission of a robustness requirement. It may stem from the complexity of political negotiations regarding the AIA, particularly regarding GPAIMs, which were not addressed in the initial draft of the regulation but gathered widespread media attention during the legislative procedure. However, evidence from ML research suggests that non-adversarial robustness is also relevant for GPAIMs (Yuan et al., 2023; Chen et al., 2022).
@@ Appendix B Notions of High-Risk AI Systems and General Purpose Models provided in AIA
An AI systems is also considered high-risk, if it operates in types or use-cases enlisted in Annex III (Art. 6(2) AIA) and poses a significant risk of harm to the health, safety or fundamental rights of natural person following (Art. 6(3) AIA). Annex III enlists eight types or use-cases of AI systems: biometric applications (e.g., remote biometric identification systems excluding biometric verification, biometric categorisation, and emotion recognition); critical infrastructure (e.g., “critical digital infrastructure, road traffic, or in the supply of water, gas, heating or electricity”); education and vocational training; employment, workers management and access to self-employment; essential private and public services; law enforcement; migration, asylum and border control management; and administration of justice and democratic processes. Thereby, the Commission shall “provide […] a comprehensive list of practical examples of use cases of AI systems that are high-risk and not high-risk” (Art. 6(5) AIA). Rec. (52) suggests that “it is appropriate to classify them [i.e., AI systems] as high-risk if, in light of their intended purpose, they pose a high risk of harm to the health and safety or the fundamental rights of persons, taking into account both the severity of the possible harm and its probability of occurrence and they are used in a number of specifically pre-defined areas”. Thereby the methodology and citeria for the identification of high-risk should be able to be adopted in order to account for “the rapid pace of technological development, as well as the potential changes in the use of AI systems” (Rec. (52)). For example, an AI system used for remote biometric identification is considered high-risk, unless it does not pose a significant risk. Remote identification refers to comparing biometric data of that individual to stored biometric data of individuals in a reference database (Rec. (15)), where a remote biometric identification system should be understood as a “AI system intended for the identification of natural persons without their active involvement, typically at a distance” (Rec. (17)).
@@
The term ’relevant circumstance’ is not defined in the AIA and therefore requires interpretation. On the one hand, one could argue that the term only refers to circumstances that are “important” for a “particular purpose” or context (Cambridge University Press, 2024c). On the other hand, the meaning of the term can also result from a comparison with other provisions of the AIA such as Art. 13(3)(b)(ii) AIA, which suggests a different understanding. The provision demands that the instructions for the use of AI systems shall contain “any known and foreseeable circumstances” that may have an impact on cybersecurity. This speaks for a broader understanding of relevance, which only excludes unknown and unforeseeable circumstances. Given this ambiguity, standards should elaborate on how to determine relevant circumstances.

### v2

@@ 5. Requirements for High-Risk AI Systems
In addition to these attacks, Art. 15(5)(iii) AIA lists ‘model flaws’ as an AI-specific vulnerability. This is a vague legal term and lacks an established counterpart in the ML literature. In software contexts, the word flaw often refers to so-called bugs, which are typically the result of human errors in the coding process (Kumar and Anderson, 2023; Nissenbaum, 1996). However, the term model flaw follows the list of attacks outlined above, which are instead designed to exploit the default properties of a properly functioning ML model, and are not directly the results of errors in the coding process. Thus, it is unclear what model flaw refers to in this context, and whether technical solutions are only expected to address traditional bugs or coding errors, or whether they should address other ways of exploiting AI-specific vulnerabilities. Given that the term is situated within the cybersecurity requirements for AI system outlined in Art. 15(5) AIA, we argue that the term model flaws should be interpreted as flaws that enable the exploitation of AI-specific vulnerabilities. Technical standards and guidelines by the EU Commission should define model flaws more clearly and provide guidelines for technical solutions to address these model flaws. This should take into account the arms race between attacker and defender in the realm of adversarial robustness, in which both parties are continuously adapting their strategies to outmanoeuvre each other (Chen et al., 2017b). This makes it infeasible to anticipate and counter all potential attacks that target AI-specific vulnerabilities.
@@ 6. Requirements for General-Purpose AI Models With Systemic Risk
In the previous section, we examined HRAIS requirements. To further elucidate them, we study GPAIMs with systemic risk, highlighting similarities and differences. The AIA establishes legal requirements for GPAIMs, such as multimodal large language models (OpenAI, 2023; Gemini Team et al., 2023), which can perform tasks beyond their original training objective (Gutierrez et al., 2023). GPAIM can be stand-alone or embedded in an HRAIS, with the latter requiring compliance with both GPAIM and HRAIS requirements. The AIA distinguishes between GPAIM with systemic risks and those without. Art. 3(65) AIA defines ‘systemic risk’ as the risk that is specific to the high-impact capabilities of GPAIMs that have a “significant impact” on the market, public health, safety, security, fundamental rights, or society. [^19] A systemic risk is presumed when the cumulative computation during training exceeds $10^{25}$ Floating-Point Operations Per Second (FLOPS). GPAIMs with fewer FLOPS may still be classified as posing a systemic risk under Art. 51(1) AIA. There is an ongoing debate over this threshold and if a model’s complexity truly reflects its risk level (Hacker, 2024; Novelli et al., 2024; Pehlivan, 2024; Kutterer, 2023). Thresholds and criteria can be modified by the EU Commission, see Appendix B GPAIMs without systemic risks are exempt from robustness and cybersecurity obligations (Art. 53 AIA ff.).
@@
Art. 55(1)(d) AIA mandates “an adequate level of cybersecurity protection” for GPAIMs with systemic risk. Rec. (115) further details this cybersecurity requirement. It mandates cybersecurity protection against “malicious use or attacks” and lists specific adversarial threats, such as “accidental model leakage, unauthorised releases, circumvention of safety measures”, “cyberattacks”, or “model theft”. Notably, several of these threats have direct counterparts in the ML literature on adversarial robustness and privacy for large generative models, such as the circumvention of safety measures (jailbreaking) or model theft (Yao et al., 2024; Li et al., 2023; Wang et al., 2023). Although Art. 55(1)(d) AIA does not define the term ‘cyberattacks’, we infer that it includes the attacks exploiting AI-specific vulnerabilities mentioned in Art. 15 AIA (see Section 5.3). These attacks are studied in the field of adversarial robustness and can also affect GPAIMs (Qiang et al., 2024; Das et al., 2024; Yan et al., 2024; Schwinn et al., 2024; Vitorino et al., 2024)—even though specific ML techniques may be necessary to address GPAIM-specific challenges. This relation underscores that the concepts and problems explored under adversarial robustness are reflected in the term ‘cybersecurity’ as used in Art. 55(1)(d) AIA. To ensure the cybersecurity of a GPAIMs with systemic risk, providers must conduct and document internal and/or external adversarial testing of the model, such as red teaming. [^20] In this context, red teaming refers to stress testing AI models by simulating adversarial attacks (Feffer et al., 2024), such as linguistic or semantic attacks against LLMs (Shi et al., 2024), whereas traditional cybersecurity red teaming focuses on assessing entire systems or networks. (Teichmann and Boticiu, 2023)
While the AIA mandates robustness requirements for HRAIS, we observe that it does not impose an explicit equivalent legal requirement for GPAIMs, regardless of whether they present a systemic risk or not. Specifically, neither Art. 55 AIA nor Rec. (155) address unintentional causes for deviations from consistent performance. In Section 5.2, we stated that non-adversarial robustness is reflected in the term robustness in Art. 15 AIA. Consequently, GPAIMs, which are not required to fulfil any robustness requirement, are not mandated to be resilient against performance issues, such as data distribution shifts or noisy data. The AIA itself does not provide an explanation for the omission of a robustness requirement. It may stem from the complexity of political negotiations regarding the AIA, particularly regarding GPAIMs, which were not addressed in the initial draft of the regulation but gathered widespread media attention during the legislative procedure. However, evidence from ML research suggests that non-adversarial robustness is also relevant for GPAIMs (Yuan et al., 2023; Chen et al., 2022).
@@ 7. Summary and Outlook
Acknowledgements.
Thank you to Tommaso Fia, and Sebastian Bordt for helpful feedback and comments. Special thank you to Marie-Sophie Müller for a thorough review of our text. Michèle Finck is a members of the Machine Learning Cluster of Excellence, funded by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645. Michèle Finck and Henrik Nolte thank the Carl Zeiss Foundation for funding support. Miriam Rateike is grateful for the generous funding support by the 2023 Google PhD Fellowship in Machine Learning.
@@ Appendix B Notions of High-Risk AI Systems and General Purpose Models provided in AIA
An AI systems is also considered high-risk, if it operates in types or use-cases enlisted in Annex III (Art. 6(2) AIA) and poses a significant risk of harm to the health, safety or fundamental rights of natural person following (Art. 6(3) AIA). Annex III enlists eight types or use-cases of AI systems: biometric applications (e.g., remote biometric identification systems excluding biometric verification, biometric categorisation, and emotion recognition); critical infrastructure (e.g., “critical digital infrastructure, road traffic, or in the supply of water, gas, heating or electricity”); education and vocational training; employment, workers management and access to self-employment; essential private and public services; law enforcement; migration, asylum and border control management; and administration of justice and democratic processes. Thereby, the Commission shall “provide […] a comprehensive list of practical examples of use cases of AI systems that are high-risk and not high-risk” (Art. 6(5) AIA). Rec. (52) suggests that “it is appropriate to classify them [i.e., AI systems] as high-risk if, in light of their intended purpose, they pose a high risk of harm to the health and safety or the fundamental rights of persons, taking into account both the severity of the possible harm and its probability of occurrence and they are used in a number of specifically pre-defined areas”. Thereby the methodology and criteria for the identification of high-risk should be able to be adopted in order to account for “the rapid pace of technological development, as well as the potential changes in the use of AI systems” (Rec. (52)). For example, an AI system used for remote biometric identification is considered high-risk, unless it does not pose a significant risk. Remote identification refers to comparing biometric data of that individual to stored biometric data of individuals in a reference database (Rec. (15)), where a remote biometric identification system should be understood as a “AI system intended for the identification of natural persons without their active involvement, typically at a distance” (Rec. (17)).
@@
The term ‘relevant circumstance’ is not defined in the AIA and therefore requires interpretation. On the one hand, one could argue that the term only refers to circumstances that are “important” for a “particular purpose” or context (Cambridge University Press, 2024c). On the other hand, the meaning of the term can also result from a comparison with other provisions of the AIA such as Art. 13(3)(b)(ii) AIA, which suggests a different understanding. The provision demands that the instructions for the use of AI systems shall contain “any known and foreseeable circumstances” that may have an impact on cybersecurity. This speaks for a broader understanding of relevance, which only excludes unknown and unforeseeable circumstances. Given this ambiguity, standards should elaborate on how to determine relevant circumstances.